OpenStudy (anonymous):

Find the radius of convergence if |3x+2| lim n->inf a_n <1, and the limit equals 1. The answer is {2/6}, but i dont understand why. See complete problem below.

OpenStudy (anonymous):

\[|3x+2| \lim_{n \rightarrow \infty} an <1\] \[|3x+2| <1\] = \[-1<3x+2< -1\] = \[-1< x < -1/3\] and some how, radius of convergence= {2/6}

OpenStudy (amistre64):

radius of convergence is the distance from the midpoint to an and point

OpenStudy (amistre64):

end point that is

OpenStudy (amistre64):

as such, radius is going to be postive in value

OpenStudy (amistre64):

now you sure its 2/6? which equals 1/3

OpenStudy (anonymous):

Thats what he said and wrote on the board. Which is also y i was a bit confused. Its as if he multiplied spomething to get that number, even though I could just simply to get 1/3

OpenStudy (anonymous):

thats what i would have done. Teachers do unnecessary things sometimes

OpenStudy (amistre64):

also, notice that the midpoint of -1 to -1/3 add and divide in half 1+1/3 = 4/3 4/3 ---- = 4/6 = 2/3 2 so the 2/6 might be a typo

OpenStudy (anonymous):

Ah, okay. Thanks!